You must evaluate a proposal to buy a new milling machine. The purchase price of the milling machine, including shipping and installation costs, is $198,000, and the equipment will be fully depreciated at the time of purchase. The machine would be sold after 3 years for $127,000. The machine would require a $3,500 increase in net operating working capital (increased inventory less increased accounts payable). There would be no effect on revenues, but pretax labor costs would decline by $54,000 per year. The marginal tax rate is 25%, and the WACC is 9%. Also, the firm spent $4,500 last year investigating the feasibility of using the machine. How should the $4,500 spent last year be handled? The cost of research is an incremental cash flow and should be included in the analysis. Only the tax effect of the research expenses should be included in the analysis. Last year's expenditure should be treated as a terminal cash flow and dealt with at the end of the project's life. Hence, it should not be included in the initial investment outlay. Last year's expenditure is considered an opportunity cost and does not represent an incremental cash flow. Hence, it should not be included in the analysis. Last year's expenditure is considered a sunk cost and does not represent an incremental cash flow. Hence, it should not be included in the analysis. What is the initial investment outlay for the machine for capital budgeting purposes after the 100% bonus depreciation is considered, that is, what is the Year 0 project cash flow? Enter your answer as a positive value. Round your answer to the nearest dollar. $ What are the project's annual cash flows during Years 1, 2, and 3? Do not round intermediate calculations. Round your answers to the nearest dollar. Year 1: $ Year 2: $ Year 3: $ Should the machine be purchased?
Answer #1

1) Last year's expenditure is considered a sunk cost and does not represent an incremental cash flow. Hence, it should not be included in the analysis.

2) You get 25% tax benefit for the bonus depreciation. So, in actual, the outflow is only 75%.

Year 0 cash flow = $198,000 x 75% + $3,500 = $152,000

3) Year 1 and 2 cash flow = Savings in labor costs x (1 - tax rate) = $54,000 x (1 - 25%) = $40,500

Year 3 cash flow = $40,500 + $3,500 + $127,000 x (1 - 25%) = $139,250

4) NPV = (-)Year 0 cash flow + Year 1 cash flow x PVIF(9%, 1) + Year 2 cash flow x PVIF(9%, 2) + Year 3 cash flow x PVIF(9%, 3)

NPV = (-)152,000 + 40,500 x 0.9174311927 + 40,500 x 0.8416799933 + 139,250 x 0.7721834801 = $26,770.55

Since NPV is positive, the machine should be purchased.